### v2

@@ Abstract
Recently, introducing Tensor Decomposition (TD) techniques into unsupervised feature selection (UFS) has been an emerging research topic. A tensor structure is beneficial for mining the relations between different modes and helps relieve the computation burden. However, while existing methods exploit TD to preserve the data tensor structure, they do not consider the influence of data orientation and thus have difficulty in handling orientation-specific data such as time series. To solve the above problem, we utilize the orientation-dependent tensor-tensor product from Tensor Singular Value Decomposition based on $\star_{\mathbf{M}}$-product (T-SVDM) and extend the one-dimensional Sparse Principal Component Analysis (SPCA) to a tensor form. The proposed sparse tensor PCA model can constrain sparsity at the specified mode and yield sparse tensor principal components, enhancing flexibility and accuracy in learning feature relations. To ensure fast convergence and a flexible description of feature correlation, we develop a convex version specially designed for general UFS tasks and propose an efficient slice-by-slice algorithm that performs dual optimization in the transform domain. Experimental results on real-world datasets demonstrate the effectiveness and remarkable computational efficiency of the proposed method for tensor data of diverse structures over the state-of-the-arts. With a proper combination of data orientation and transform domain, our method is promising for various applications. The codes related to our proposed methods and the experiments are available at https://github.com/zjj20212035/STPCA.git.
@@ 1 Introduction
Feature Selection (FS) is a technique of selecting as few features (variables) as possible to meet the demands of a certain task [1, 2]. It is usually employed as a pre-processing algorithm to relieve the computational burden and enhance interpretability by offering only the necessary features. According to the amount of supervised information during training, FS methods can be classified as supervised [3], semi-supervised [4], and unsupervised [5]. Among them, unsupervised feature selection (UFS) is more challenging due to the unavailability of labels. To overcome the difficulty, most UFS methods adopt learning methods including but not limited to principal component analysis (PCA) [38], spectral clustering [40], and graph-based learning [18]. In this paper, we focus on PCA-based UFS methods.
In most UFS methods, data samples are vectorized, resulting in the curse of dimensionality [6] as well as data structure loss when dealing with tensors. As a natural extension of matrix to store multi-way data samples, tensor is now a common processed object in various communities such as image processing [7], signal processing [8], and economical analysis[9]. A $n$-th order tensor comprises $n$ modes, each containing several dimensions. The tensor structure contains rich information on the relations of modes and dimensions, which is beneficial for analyzing the features. In this paper, we focus on leveraging the tensor structure to reduce computational complexity without compromising UFS performance.
To utilize the structure information, tensor-based UFS methods apply all kinds of Tensor Decomposition (TD) techniques[18, 19, 20]. Tensor decomposition aims to factorize a tensor into a product of several subtensors, which has been considerably studied for decades since the famous Tucker decomposition [21] was proposed in 1964. As different definitions for tensor rank are established, various TD methods have been put forward, such as Higher-order Singular Value Decomposition (HOSVD) [12], CANDECOMP/PARAFAC decomposition (CPD) [22, 23], and Tensor Singular Value Decomposition (T-SVD)[24]. One purpose of TD is to mine the relations between modes and dimensions, which are naturally related to UFS. TD methods preserve the tensor structure of UFS, while UFS helps validate the accuracy and interpretability of the relationships learned by the TD method. By minimizing the reconstruction error between a constructed tensor and its decomposition, tensor-based UFS methods successfully preserve the data structure information. However, most of these methods do not develop sparse TD and thus use the classifier matrix to perform UFS, which is not beneficial for interpretability and utilization of the learned element relations. To address this problem, we start by finding the compact representations of the factor matrices. As is widely known, TD methods are also viewed as ways to achieve Tensor Principal Component Analysis (TPCA) [12, 25]. The principal components have the greatest variance among data samples on a global manifold. Their directions describe the feature necessity during the projection onto the manifold. Therefore, the key to selecting interpretable and discriminative features could be hidden in TPCA.
In terms of performing linear projection, the current TPCA methods can be roughly divided into two types: Tuckers-based TPCA [12, 26, 27] and T-SVDM-based TPCA [28, 29, 30]. Tuckers-based TPCA applies direction-unfolding products defined by Tucker decomposition and its variations (including CPD, which can be viewed as a special case of Tucker’s), while T-SVDM-based TPCA is based on the $\star_{\mathbf{M}}$-product defined by T-SVD and its generalization T-SVDM [31]. The main differences between the above two types of methods lie in the following two aspects: 1) Tucker-based methods conduct tensor-matrix products through tensor matricization in the original domain, using factor matrices, while T-SVDM-based methods perform tensor-tensor products in the transform domain with a factor tensor. By applying factor tensors, T-SVDM-based methods obtain significantly more weights assigned for elements in the data tensor, describing feature relations in a more delicate way. 2) Tucker-based methods obtain identical results regardless of data organization, while T-SVDM-based methods can achieve various decompositions by applying different transform domains and changing tensor orientation. The reorientation works as prior knowledge of mode relations and guides the T-SVDM-based methods to further learn the feature relations in a proper direction, especially in some orientation-specific scenarios such as time series. Moreover, the transform operation provides T-SVDM-based methods with greater flexibility, allowing their application to various scenarios. In a word, T-SVDM-based methods have a unique advantage in handling complex scenarios. We are interested in introducing T-SVDM-based TPCA into UFS. To achieve this goal, an extension to a sparse framework is required to ensure the preservation of only the necessary feature relations. To the best of our knowledge, this framework is still absent.
@@
This paper develops a novel sparse T-SVDM-based TPCA framework and deduces a convex version for UFS. We explore a way of leveraging the tensor structure to reduce computational complexity while preserving satisfying UFS performance. Our contributions are summarized as follows:
@@ 2 Related Work
In this section, we introduce notations used throughout the paper, and research revolving around tensor decomposition, tensor PCA, and UFS.
@@ 2.1.1 Basic Symbols
In most cases, we use capital letters (e.g. $A$) to denote sets, except the common ones like integer set $\mathbb{Z}$, the real-number set $\mathbb{R}$ and the complex-number set $\mathbb{C}$. In particular, $H^{d}_{+}$ stands for the HPSD Cone for $d\times{d}$ matrices, and $\Phi$ is the empty set. For an integer $n$, $[n]$ represents the set $\left[1,2,3,...,n\right]$. #$\#(A)$ stands for the number of elements in the set $A$. A set with the superscript "o" has ordered elements; for example, $A^{o}=\{1,2,3\}\neq B^{o}=\{1,3,2\}$.
We use lowercase letters (e.g. $a$) to denote scalars, boldface lowercase letters (e.g. $\mathbf{a}$) to denote vectors, boldface capital letters (e.g. $\mathbf{A}$) and Euler script letters (e.g. $\mathcal{A}$) to denote tensors. For a $n^{th}$-order tensor $\mathcal{A}\in\mathbb{C}^{d_{1}\times{d_{2}}\times{\cdots}\times{d_{n}}}$, its $(i_{1},i_{2},\cdots,i_{n})$-th element is denoted by $\mathcal{A}_{i_{1}i_{2}\cdots i_{n}}$. For a matrix, the corresponding element is denoted by $a_{i_{1}i_{2}}$. For a matrix $\mathbf{A}$, its row vectors are denoted by $\mathbf{A}_{j,:}$ or boldface lowercase letters $\mathbf{a}^{j}$. The column vectors are denoted by $\mathbf{A}_{:,j}$ or boldface lowercase letters $\mathbf{a}_{j}$.
There are two types of common sections for third-order tensors: fibers and slices. The vector along the $n$-th mode is referred to as an $n$-mode fiber. A third-order tensor $\mathcal{A}\in\mathbb{C}^{d_{1}\times{d_{2}}\times{d_{3}}}$ has column (1-mode) fibers $\mathcal{A}_{:i_{2}i_{3}}$, row (2-mode) fibers $\mathcal{A}_{i_{1}:i_{3}}$, and tube (3-mode) fibers $\mathcal{A}_{i_{1}i_{2}:}$. A slice is a matrix obtained by fixing all but two indices in a tensor. For a third-order tensor, slices can be divided into frontal slices $\mathcal{A}_{::i_{3}}$, lateral slices $\mathcal{A}_{:i_{2}:}$ and horizontal slices $\mathcal{A}_{i_{1}::}$.
The transpose, conjugation, and conjugate transpose of a matrix are denoted by $\mathbf{A}^{T}$, $\mathbf{A}^{*}$, and $\mathbf{A}^{H}$, respectively. We denote the trace of a square matrix by ${\rm Tr}(\mathbf{A})$. We denote Frobenius norm, nuclear norm and $\ell_{2,1}$ norm by ${\|\mathcal{A}\|}_{F}$,${\|\mathbf{A}\|}_{*}$, and ${\|\mathbf{A}\|}_{2,1}$ respectively. The nuclear norm of a matrix is defined as the sum of its singular values. The $\ell_{2,1}$-norm is obtained by calculating the $\ell_{2}$-norm of each column vector in a matrix and summing the results. The Kronecker product is denoted by $\bigotimes$.
@@ 2.1.2 Tensor Operations
In this section, we introduce several tensor operations.
@@
We use $n$-mode unfolding to perform tensor matricization. A $m$-th order tensor $\mathcal{A}^{d_{1}\times{d_{2}}\dots\times{d_{m}}}$ can be transformed into $\mathbf{X}_{(n)}$ along the $n$-th mode where
@@
Using the unfolding technique, the tensor-matrix product is defined as
@@
In the T-SVDM framework, third-order tensor products are achieved through the $\star_{\mathbf{M}}$-product. Given a third-order tensor $\mathcal{A}\in\mathbb{C}^{d\times{m}\times{n}}$, we first denote a tensor in the transform domain specified by an invertible matrix $\mathbf{M}\in\mathbb{C}^{n\times{n}}$:
@@
Then, given another tensor $\mathcal{B}\in\mathbb{C}^{m\times{l}\times{n}}$, $\mathcal{A}\star_{\mathbf{M}}\mathcal{B}=\mathcal{C}$ is defined as
$\hat{\mathcal{C}}_{::i}$ $=\hat{\mathcal{A}}_{::i}\hat{\mathcal{B}}_{::i},\quad i=1,2,\cdots,n,$ (4)
@@
Hereafter, we use $\hat{\mathcal{A}}$ to denote the result of $\mathcal{A}\times_{3}\mathbf{M}$. The $\star_{\mathbf{M}}$-product is a slice-by-slice operation and is thus orientation-dependent. Different orientations of the tensor will lead to different results.
The identity tensor $\mathcal{I}$ under $\star_{\mathbf{M}}$ satisfies the following condition:
@@
Each frontal slice of $\hat{\mathcal{I}}$ is an identity matrix $\mathbf{I}$.
@@
$\left(\hat{\mathcal{A}}^{H}\right)_{::i}=\left(\hat{\mathcal{A}}_{::i}\right)^{H},\quad i=1,\cdots,n,$ (6)
which makes sure that $\mathcal{A}^{H}\star_{\mathbf{M}}\mathcal{B}^{H}=\left(\mathcal{B}\star_{\mathbf{M}}\mathcal{A}\right)^{H}$. We can further define a $\star_{\mathbf{M}}$-orthogonal tensor $\mathcal{A}\in{\mathbb{C}^{m\times{m}\times{n}}}$ if
@@ 2.2 T-SVDM
Using the $\star_{\mathbf{M}}$ product, T-SVDM [31] decomposes a third-order tensor $\mathcal{A}\in\mathbb{C}^{d_{1}\times{d_{2}}\times{d_{3}}}$ into:
@@
where $\mathcal{U}\in\mathbb{C}^{d_{1}\times d_{1}\times d_{3}}$, $\mathcal{V}\in\mathbb{C}^{d_{2}\times d_{2}\times d_{3}}$ are $\star_{\mathbf{M}}-unitary$, and $\mathcal{S}\in\mathbb{C}^{d_{1}\times{d_{2}}\times{d_{3}}}$ is a f-diagonal tensor whose frontal slices are diagonal. The decomposition is called T-SVD [24] when $\mathbf{M}$ is the discrete Fourier transform (DFT) matrix and T-SVDM [31] in a usual case.
[28] was the first to establish the relationship between T-SVD and tensor PCA. It applies T-SVD to compress the human-face images with a pre-rotation of the data tensor. [31] further extends this idea to T-SVDM. [29] highlights the importance of leveraging the orientation-dependence of T-SVDs to extract information from different directions. Inspired by these works, we derive a sparse model for tensors combining T-SVDM and SPCA.

### v3

@@ Abstract
Recently, introducing Tensor Decomposition (TD) techniques into unsupervised feature selection (UFS) has been an emerging research topic. A tensor structure is beneficial for mining the relations between different modes and helps relieve the computation burden. However, while existing methods exploit TD to preserve the data tensor structure, they do not consider the influence of data orientation and thus have difficulty in handling orientation-specific data such as time series. To solve the above problem, we utilize the orientation-dependent tensor-tensor product from Tensor Singular Value Decomposition based on $\star_{\mathbf{M}}$-product (T-SVDM) and extend the one-dimensional Sparse Principal Component Analysis (SPCA) to a tensor form. The proposed sparse tensor PCA model can constrain sparsity at the specified mode and yield sparse tensor principal components, which enhances flexibility and accuracy in learning feature relations. To ensure fast convergence and a flexible description of feature correlation, we develop a convex version specially designed for general UFS tasks and propose an efficient slice-by-slice algorithm that performs dual optimization in the transform domain. Experimental results on real-world datasets demonstrate the effectiveness and remarkable computational efficiency of the proposed method for tensor data of diverse structures over the state-of-the-art. When transform axes align with feature distribution patterns, our method is promising for various applications. The codes related to our proposed method and the experiments are available at https://github.com/zjj20212035/STPCA.git.
@@ 1 Introduction
Feature selection (FS) aims to select a minimal subset of features (variables) sufficient for a given task [1, 2]. Typically employed as a pre-processing step, FS reduces computational complexity and enhances interpretability by providing only essential features. FS methods are commonly categorized based on the level of supervision during training: supervised [3], semi-supervised [4], and unsupervised [5]. Among these, Unsupervised Feature Selection (UFS) presents a greater challenge due to the absence of labels. To address this, most UFS methods leverage learning techniques, including but not limited to principal component analysis (PCA) [41], spectral clustering [43], and graph-based learning [21]. This paper specifically focuses on PCA-based UFS methods.
A common limitation in many UFS approaches is the vectorization of data samples, which can lead to the curse of dimensionality [6] and loss of inherent data structure, particularly when handling tensors. Tensors, as natural multi-way extensions of matrices for representing data, are prevalent in diverse fields like image processing [7], signal processing [8], and econometric analysis [9]. An $n$-th order tensor consists of $n$ modes, each with multiple dimensions, inherently encoding rich relational information between these modes and dimensions that is valuable for feature analysis. This paper focuses on exploiting this tensor structure to reduce computational complexity without compromising UFS performance. performance.
To leverage structural information, tensor-based UFS methods employ various Tensor Decomposition (TD) techniques [21, 22, 23]. Tensor decomposition factorizes a tensor into a product of several subtensors. This area has been extensively studied since the seminal Tucker decomposition [24] was introduced in 1964. The establishment of different tensor rank definitions has led to the development of diverse TD methods, such as Higher-Order Singular Value Decomposition (HOSVD) [12], CANDECOMP/PARAFAC decomposition (CPD) [25, 26], and Tensor Singular Value Decomposition (T-SVD) [27].
A key purpose of TD is to uncover the relationships between modes and dimensions, which are intrinsically linked to UFS. TD methods preserve the tensor structure crucial for UFS, while UFS provides a framework to validate the accuracy and interpretability of the relationships learned through TD. By minimizing the reconstruction error between the original tensor and its decomposed representation, tensor-based UFS methods effectively preserve data structure information. However, most existing methods do not inherently impose sparsity constraints within the TD process itself. Consequently, they often rely on a separate classifier matrix to perform feature selection, which can obscure the interpretability and limit the direct utilization of the learned element relationships. To address this limitation, we propose seeking compact representations of the factor matrices. It is well-established that TD methods can also be interpreted as achieving Tensor Principal Component Analysis (TPCA) [12, 28]. Principal components capture maximal data variance, and their directions indicate the significance of features during projection onto the underlying manifold. Therefore, the key to selecting interpretable and discriminative features likely lies in TPCA.
Recent studies in domains such as electroencephalogram (EEG) analysis [17] underscore the necessity of orientation-aware approaches. [15] demonstrates that spatial-only feature representations, which neglect time-frequency couplings, fail to capture discriminative features in multi-channel EEG signals with temporal variability, resulting in 15%-26% silhouette score degradation for feature selection and clustering tasks. Consequently, the authors implement feature selection separately across time, frequency, and time-frequency domains, significantly enhancing clustering accuracy. In broader applications, TPCA methods can be categorized into three primary types: Tucker-based TPCA [12, 29, 30], CPD-based TPCA, and T-SVDM-based TPCA [31, 32, 33]. Key distinctions are summarized in Table 1. While Tucker-based and CPD-based methods model potential mode correlations, their reliance on vector outer product principles renders them inherently orientation-insensitive [27]. This limitation manifests in orientation-specific data (e.g., temporal sequences), where Tucker’s and CPD’s decomposition results remain invariant to data orientation. In contrast, T-SVDM-based approaches leverage matrix outer products to adapt to orientation and transform domain variations, yielding direction-sensitive outcomes with broader applicability. Furthermore, T-SVDM’s slice-by-slice processing (Section 2.1.2) enables slice-wise sparsity, allowing independent optimization of factor matrices. This offers greater flexibility for algorithmic design compared to the rank-constrained CPD and Tucker decomposition that produces dense core tensors. [34] further highlights T-SVDM’s reduced parameter sensitivity versus Tucker and CPD in image denoising. Although T-SVDM-based methods incur higher computational complexity from factor tensor operations, they provide granular element-wise weighting that meticulously characterizes feature relationships. These attributes confer unique advantages for complex scenarios. Motivated by these strengths, we propose integrating T-SVDM-based TPCA into UFS. Achieving this requires extending the framework with sparsity constraints to preserve only essential feature relations. To our knowledge, such a framework remains absent in the literature.
Table 1: Comparative analysis of UFS methods based on different TD approaches.
Property Tucker-based CPD-based T-SVDM-based
Basic product vector outer product vector outer product matrix outer product
Computational Complexity $O(t(d_{1}d_{2}k_{1}k_{2}+d_{1}d_{3}k_{1}k_{3}+d_{2}d_{3}k_{2}k_{3}))^{\ast}$ $O(t(d_{1}d_{2}+d_{1}d_{3}+d_{2}d_{3})k^{2})$ $O\big{(}d_{3}^{2}d_{1}d_{2}+{}$
$d_{3}\min(d_{1}d_{2}^{2},d_{1}^{2}d_{2})\big{)}$
Sensitivity to Orientation Isotropic Permutation-invariant Directionally aware
Sparsity Handling Dense core Rank-constrained Slice-wise sparsity
Domain Adaptability Original domain Original domain Transform domain
$t$: iterations. $d_{i}$: mode-$i$ dimensionality (original tensor). $k_{i}$: mode-$i$ dimensionality (core tensor) ($i=1,2,3$).
@@
This paper proposes a novel sparse T-SVDM-based TPCA framework with a convex formulation for Unsupervised Feature Selection. We establish a methodology that leverages tensor structures to reduce computational complexity while maintaining competitive UFS performance. Our contributions are summarized as follows:
@@ 2 Related Work
This section introduces mathematical notations employed throughout the paper and reviews relevant research on tensor decomposition, tensor principal component analysis, and unsupervised feature selection.
@@ 2.1.1 Basic Symbols
Notations employed throughout the paper are summarized in Table 2. Specifically, fibers correspond to vector sections along individual tensor modes, while slices represent matrix sections defined by fixing all but two modes.
Table 2: Summary of mathematical notations. The following notations use the letter A as an example, and they are equally applicable to other letters.
Symbol Description
$A$ Generic set
$\mathbb{Z},\mathbb{R},\mathbb{C}$ Integer, real, complex number sets
$H^{d}_{+}$ Hermitian Positive Semidefinite (HPSD) cone ($d\times d$ matrices)
$\Phi$ Empty set
$[n]$ Ordered set $\{1,2,\dots,n\}$
$A^{o}$ Ordered set (element order matters)
#$\#(A)$ Cardinality of set $A$
$a$ Scalar
$\mathbf{a}$ Vector
$\mathbf{A}$ Matrix
$\mathcal{A}$ Third-order Tensor
$\mathcal{A}_{:i_{2}i_{3}}$, $\mathcal{A}_{i_{1}:i_{3}}$, $\mathcal{A}_{i_{1}i_{2}:}$ Column (mode-1), Row (mode-2), and Tube (mode-3) fibers
$\mathcal{A}_{::i_{3}}$, $\mathcal{A}_{:i_{2}:}$, $\mathcal{A}_{i_{1}::}$ Frontal, Lateral and Horizontal slices
$\mathcal{A}_{i_{1}\cdots i_{n}}$ Tensor element at $(i_{1},\dots,i_{n})$
$a_{ij}$ Matrix element at $(i,j)$
$\mathbf{A}_{j,:},\mathbf{a}^{j}$ $j$-th row vector of matrix
$\mathbf{A}_{:,j},\mathbf{a}_{j}$ $j$-th column vector of matrix
$\mathbf{A}^{T},\mathbf{A}^{*},\mathbf{A}^{H}$ Transpose, conjugate, conjugate transpose
$\mathrm{Tr}(\mathbf{A})$ Matrix trace
$\bigotimes$ Kronecker product
$\|\mathcal{A}\|_{F}$ Frobenius norm ($\sqrt{\sum_{i_{1}}\sum_{i_{2}}\sum_{i_{3}}\mathcal{A}^{2}_{i_{1}i_{2}i_{3}}}$)
$\|\mathbf{A}\|_{*}$ Nuclear norm ($\sum\sigma_{i}^{\ast}$)
$\|\mathbf{A}\|_{2,1}$ $\ell_{2,1}$ norm ($\sum_{j}\|\mathbf{A}_{:,j}\|_{2}$)
singular values
@@ 2.1.2 Tensor Operations
In this section, we introduce fundamental tensor operations.
@@
Tensor unfolding (matricization) is performed via $n$-mode unfolding. For an $m$-th order tensor $\mathcal{X}\in\mathbb{R}^{d_{1}\times d_{2}\times\cdots\times d_{m}}$, its $n$-mode unfolding $\mathbf{X}_{(n)}$ is defined such that:
@@
Based on this unfolding operation, the tensor-matrix product ($n$-mode product) is defined as:
@@
Within the T-SVDM framework, the $\star_{\mathbf{M}}$-product defines tensor operations for third-order tensors. Consider $\mathcal{A}\in\mathbb{C}^{d\times{m}\times{n}}$ and an invertible transformation matrix $\mathbf{M}\in\mathbb{C}^{n\times{n}}$. We first define the transformed tensor in the $\mathbf{M}$-domain as:
@@
For another tensor $\mathcal{B}\in\mathbb{C}^{m\times{l}\times{n}}$, the $\star_{\mathbf{M}}$-product $\mathcal{C}=\mathcal{A}\star_{\mathbf{M}}\mathcal{B}$ is given by:
$\hat{\mathcal{C}}_{::i}$ $=\hat{\mathcal{A}}_{::i}\hat{\mathcal{B}}_{::i},\quad\forall i\in\{1,2,\cdots,n\},$ (4)
@@
Throughout this work, $\hat{\mathcal{A}}$ consistently denotes $\mathcal{A}\times_{3}\mathbf{M}$. The $\star_{\mathbf{M}}$-product performs slice-by-slice matrix multiplications to achieve tensor-tensor products in the transform domain, making it inherently orientation-sensitive. Consequently, different tensor orientations yield distinct results.
The identity tensor $\mathcal{I}$ under the $\star_{\mathbf{M}}$-product satisfies:
@@
In the transform domain, each frontal slice of $\hat{\mathcal{I}}$ corresponds to an identity matrix $\mathbf{I}$.
@@
$\left(\hat{\mathcal{A}}^{H}\right)_{::i}=\left(\hat{\mathcal{A}}_{::i}\right)^{H},\quad\forall i\in\{1,\cdots,n\},$ (6)
which ensures the fundamental property:
$\mathcal{A}^{H}\star_{\mathbf{M}}\mathcal{B}^{H}=\left(\mathcal{B}\star_{\mathbf{M}}\mathcal{A}\right)^{H}.$ (7)
A tensor $\mathcal{A}\in\mathbb{C}^{m\times{m}\times{n}}$ is $\star_{\mathbf{M}}$-orthogonal if it satisfies:
@@ 2.2 T-SVDM
Within the $\star_{\mathbf{M}}$-product framework, the T-SVDM decomposition [34] factorizes a third-order tensor $\mathcal{A}\in\mathbb{C}^{d_{1}\times d_{2}\times d_{3}}$ as:
@@
where $\mathcal{U}\in\mathbb{C}^{d_{1}\times d_{1}\times d_{3}}$ and $\mathcal{V}\in\mathbb{C}^{d_{2}\times d_{2}\times d_{3}}$ are $\star_{\mathbf{M}}$-unitary tensors. $\mathcal{S}\in\mathbb{C}^{d_{1}\times d_{2}\times d_{3}}$ is an f-diagonal tensor with diagonal frontal slices. This decomposition generalizes two important cases: T-SVD [27]: Special case when $\mathbf{M}$ is the DFT matrix; T-SVDM [34]: General case with arbitrary invertible $\mathbf{M}$.
The connection between T-SVDM and TPCA was first established in [31], which applied T-SVD to facial image compression through data tensor pre-rotation. This work was later extended to T-SVDM in [34]. Notably, [32] demonstrated the critical role of orientation-dependence in T-SVD for multi-directional feature extraction. Building upon these foundations, we develop a novel sparse tensor model that integrates T-SVDM with sparse PCA.